OpenStudy (anonymous):

Find an equation in standard form for the line described... Passing through the origin and with no slope. MEDAL QUESTION

OpenStudy (anonymous):

whats the line described..?

OpenStudy (anonymous):

The line described is : Passing through the origin and with no slope.

OpenStudy (anonymous):

ohhh.. sorry my bad.. uhhhh...

OpenStudy (anonymous):

x=0

OpenStudy (anonymous):

how'd u get that? can u explain?

OpenStudy (anonymous):

a vertical line has an undefined slope (no slope).. so the line has to be vertical running through the origin (0,0) which makes it running right along the Y-axis... which is x=0

OpenStudy (charlotte123):

but wouldn't it be y = 0? :P

OpenStudy (charlotte123):

x = 0 and y = 0 xD

OpenStudy (anonymous):

no.. .-. because is Y is zero, x can be anything.. which would make it horizontal... and if you have both, its just a point 0,0

OpenStudy (anonymous):

with x=0, Y can be anything.. but X has to be zero... which is... vertical .-.

OpenStudy (charlotte123):

this site made meh confused .-. http://forums.welltrainedmind.com/topic/487984-math-help-finding-equation-in-standard-form-for-line-with-no-slope/ sorry michael xD

OpenStudy (anonymous):

its okay xD

OpenStudy (charlotte123):

people were having the same problem, trying to figure out if it was y = 0 or x = 0 xD but u r right, it is x = 0! the equation has to be x= #. You're going through the origin, so x=0. as one says ^_^

OpenStudy (charlotte123):

OH WAIT! y = 0 is for the line, but x = 0 IS FOR THE EQUATION! :D

OpenStudy (charlotte123):

If "no slope" means "zero slope", the line would be y=0: y=mx+c m=0 must go through origin, so y-intercept is c=0 y=0 If "no slope" means slope goes to undefined, the line would be x=0. (I would consider the problem poorly worded.). m=(y2-y1)/(x2-x1) is undefined if denominator is zero, i.e. x2=x1. Since x=0 must be included, x=0 is the equation.

OpenStudy (anonymous):

what O.o x=0 is the equation xD

OpenStudy (charlotte123):

yep ^_^

OpenStudy (anonymous):

no slope isn't a 0 slope.. a 0 slope is still a slope.. no slope means undefined..

OpenStudy (charlotte123):

I know michael xD

OpenStudy (charlotte123):

don't say dat to me, say it to dat guy on der xD either way, x = 0 is the equation ^_^

OpenStudy (anonymous):

lol yes

OpenStudy (charlotte123):

I am done wid dis confusion .-. XD

OpenStudy (anonymous):

how do i show that though?

OpenStudy (anonymous):

I got y-0=0(x-0) y=0

OpenStudy (anonymous):

uhh... i don't know if theres work to show for it.. its just kinda known that y=0 is a horizontal like and x=0 is a vertical one

OpenStudy (anonymous):

can a horizontal line have a slope?

OpenStudy (anonymous):

yes. a horizontal line has a slope of zero. but a vertical one is undefined. picture a little guy running on the line. he can run along the flat one but theres nothing to run on with the vertical one o.o

OpenStudy (anonymous):

that's a horrible analogy XD but yeah.. the first part.. ^ lol

OpenStudy (anonymous):

ohhhh i think i get it now.